1) (x-1)^2/9+(y-3)^2/4=1
what is the center?Vertices?Foci? and Co Vertices. In short give the conic section.

Answer by lwsshak3(11628)

(x-1)^2/9+(y-3)^2/4=1
what is the center?Vertices?Foci? and Co Vertices.
...
(x-1)^2/9+(y-3)^2/4=1
Standard form for ellipse with horizontal major axis: (x-h)^2/a^2+(y-k)^2/b^2=1, (a>b), with (h,k) being the (x,y) coordinates of the center.
Given ellipse has a horizontal major axis:
Center: (1,3)
a^2=9
a=3
length of major axis=2a=6
vertices: (1±3,3) or (4,3) and (-2,3)
..
b^2=4
b=2
length of co-vertices or minor axis=2b=4
end-points of minor axis=(1,3±2)
..
c^2=a^2-b^2=9-4=5
c=√5
Foci: (1±√5,3)
see graph below as a visual check on answers
..
y=±(4-4(x-1)^2/9)^.5+3